# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ 5.3.1. Concurrent $k$-sparse Parallel Composition of IM
Consider a setting where an adversary creates (an unlimited number of) IMs over time while adaptively asking queries from the existing ones. When creating an IM $\mathcal{M}_{i}$ that is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a neighbor relation $\sim_{i}$, the adversary also selects a pair of neighboring initial states for $\mathcal{M}_{i}$. This selection depends on the history of queries asked so far and their responses. Based on a secret bit $b$, either all created mechanisms are initialized with the first initial state in their respective pairs, or all are initialized with the second. The goal of the adversary is to guess $b$.
This scenario is an special case of FPC concurrent $k$-sparse parallel composition of CMs, where, instead of creating $\mathcal{M}_{i}$ and assigning a pair of initial states to it, the adversary sends $(\mathcal{M}_{i}^{\prime},\text{null},\epsilon_{i},\delta_{i},f_{\sim_{i}})^{2}$, where $\mathcal{M}_{i}^{\prime}$ is the CM simulating $\mathcal{M}_{i}$. Moreover, to ask a query $q$ from the $i$-th created IM, the adversary sends $\left((q,i),(q,i)\right)$. Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.15 applies to the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ (that are differentially private w.r.t. neighbor relations).
NIMs are a special subclass of IMs that halt the communication upon receiving a second input message. Since Theorem 5.15 applies to all IMs, it also holds for the concurrent $k$-sparse parallel composition of NIMs with privacy parameters $\left((\epsilon_{1},\delta_{i}\right)_{i=1}^{k}$. We note that the restricted version of Theorem 5.15 for the $k$-sparse parallel composition of (unbounded number of) NIMs was previously proved in [20] as an intermediate lemma. In the next section, we compare the $k$-sparse parallel composition of IMs to the classical parallel composition theorem for NIMs.
@@ 6. Designing Complex Continual Mechanisms
The privacy analysis of such mechanisms $\mathcal{M}$ is challenging as the creation of sub-mechanisms and their inputs is adaptive and dependent. The goal of this section is to give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be computed by applying the known concurrent composition theorems for continual mechanisms to the sub-mechanisms. Specifically, given that the $f^{\prime}$-concurrent composition of CMs is $(\epsilon,\delta)$-DP w.r.t. some verification function $f^{\prime}$, we would like to show that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the given verification function $f$.
@@
To prove that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. $f$, we need to show that the view of every adversary $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable. Intuitively, the messages that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ receives from $\mathcal{I}(b)$ contains information about $b$. Therefore, to deduce that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f$ from the fact that ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$, the IPM $\mathcal{P}$ must not directly use the messages from $\mathcal{I}$ to respond. Instead, it should only use these messages to determine which queries it wants to send to its its sub-mechanisms via ExtConComp. More precisely, $\mathcal{P}$ must generate an answer for $\mathcal{I}$ only based on the privatized answers provided by ExtConComp and not dependent on the un-privatized input sequence (Response Property). Moreover, $\mathcal{P}$ must decide whether to continue interacting with ExtConComp or to return a response based on these privatized answers (Destination Property).
@@
In Section 6.1, we formalize the above properties for $\mathcal{P}$, and in Section 6.2, we prove the desired privacy guarantees for $\mathcal{M}$. Combining this result with the concurrent composition theorems provided in the previous sections, we can avoid complicated privacy analysis for many CMs. To obtain rigorous results, our statements and proofs must be presented in a technical and somewhat complex form. In contrast, mechanisms are usually described more simply using pseudocode rather than message-passing state machines. In Section 6.3, we reformulate our results in terms of pseudocode, enabling new mechanisms to be analyzed directly without the need for reformulating to the setting of in this paper.
@@
Recall that the goal of this section is to show $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$ if ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f$. That is, we need to show that the view of every adversary interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
The messages received by $\mathcal{P}$ from $\mathcal{I}$ contain information about the bit used by $\mathcal{I}$ to filter the adversary’s pairs. Therefore, to directly use the existing compositions theorems for ExtConComp to analyze the privacy of $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$, we need to ensure that $\mathcal{P}$ only utilizes the messages from $\mathcal{I}$ to generate queries for the existing private sub-mechanisms executed by ExtConComp. Moreover, we need to guarantee that the sequence $(m_{1}^{0\prime},m_{1}^{1\prime}),(m_{2}^{0\prime},m_{2}^{1\prime}),\dots$ is $f^{\prime}$-valid, where $m_{1}^{b\prime},m_{2}^{b\prime},\dots$ is the sequence of messages that $\mathcal{P}$ sends to ExtConComp when it receives messages from $\mathcal{I}(b)\!\circ^{*}\!\mathcal{V}[f]$. The following properties formalize these:
@@
  • $f\to f^{\prime}$ Property. For $v\in\{{\texttt{L}},{\texttt{R}}\}$, let the pair $(m,v)$ denote a message $m$ sent to or received from the left or right mechanism of $\mathcal{P}$. For $b\in\{0,1\}$, let $\sigma^{b}=\left((m_{1}^{b},v_{1}^{b}),\dots,(m_{t}^{b},v_{t}^{b})\right)$ be a sequence of $t$ input messages to $\mathcal{P}$, and let $\tau^{b}=\left((m_{1}^{b\prime},v_{1}^{b\prime}),\dots,(m_{t}^{b\prime},v_{t}^{b\prime})\right)$ be the corresponding sequence of $t$ output messages from $\mathcal{P}$. Suppose the following conditions hold:
@@
In Theorem 6.1, we assume that the view of every adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable. The goal is to show the same result for $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$. The key idea of the proof is to show that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{T}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical, where $\mathcal{T}$ is an IPM with a singleton initial state space. The proof consists of two steps.
@@
When $\mathcal{P}^{2}$ receives a message $m\in Q_{\mathcal{P}}^{L}$ from the left mechanism, it forwards two copies of $m$ to both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ as left messages. Specifically, it executes $(s_{0},v_{0},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{L}},m)$ and $(s_{1},v_{1},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{L}},m)$. When $\mathcal{P}^{2}$ receives a pair of messages $(m_{0},m_{1})\in Q_{\mathcal{P}}^{R}\times Q_{\mathcal{P}}^{R}$ from the right mechanism, it passes each $m_{b}$ to $\mathcal{P}_{b}$ as a right message. Specifically, it executes $(s_{0},v_{0},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{R}},m_{0})$ and $(s_{1},v_{1},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{R}},m_{0})$. By the destination property of $\mathcal{P}$ and the fact that both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ receive identical left messages, we have $v_{0}^{\prime}=v_{1}^{\prime}$. If $v_{0}^{\prime}=v_{1}^{\prime}={\texttt{R}}$, then $\mathcal{P}^{2}$ sends $(m_{0}^{\prime},m_{1}^{\prime})$ to its right mechanism. If $v_{0}^{\prime}=v_{1}^{\prime}={\texttt{L}}$, then by the response property, $m_{0}^{\prime}=m_{1}^{\prime}$, and $\mathcal{P}^{2}$ returns $m_{0}^{\prime}$ to its left mechanism.
The following claim follow directly from the definition of $\mathcal{P}^{2}$:
@@
$\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}$ in Claim 6.4 is an IPM with a singleton initial state space. Thus, by Lemma 2.16, that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable implies that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof. ∎

### v4

@@ 5.3.1. Concurrent $k$-sparse Parallel Composition of IM
Consider a setting where an adversary creates (an unlimited number of) IMs over time while adaptively asking queries from the existing ones. When creating an IM $\mathcal{M}_{i}$ that is $(\epsilon_{i},\delta_{i})$-DP w.r.t. a neighbor relation $\sim_{i}$, the adversary also selects a pair of neighboring initial states for $\mathcal{M}_{i}$. This selection depends on the history of queries asked so far and their responses. Based on a secret bit $b$, either all created mechanisms are initialized with the first initial state in their respective pairs, or all are initialized with the second. The goal of the adversary is to distinguish whether $b=0$ or $b=1$.
This scenario is an special case of FPC concurrent $k$-sparse parallel composition of CMs, where, instead of creating $\mathcal{M}_{i}$ and assigning a pair of initial states to it, the adversary sends $(\mathcal{M}_{i}^{\prime},\text{null},\epsilon_{i},\delta_{i},f_{\sim_{i}})^{2}$, where $\mathcal{M}_{i}^{\prime}$ is the CM simulating $\mathcal{M}_{i}$. Moreover, to ask a query $q$ from the $i$-th created IM, the adversary sends $\left((q,i),(q,i)\right)$. Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.14 applies to the concurrent $k$-sparse parallel composition of IM (that are differentially private w.r.t. neighbor relations).
Corollary 5.15.
For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ is also $(\epsilon,\delta)$-DP.
NIMs are a special subclass of IMs that halt the communication upon receiving a second input message. Since Theorem 5.14 applies to all IMs, it also holds for the concurrent $k$-sparse parallel composition of NIMs with privacy parameters $\left((\epsilon_{1},\delta_{i}\right)_{i=1}^{k}$. In the next section, we compare the $k$-sparse parallel composition of IMs to the classical parallel composition theorem for NIMs.
@@ 6. Designing Complex Continual Mechanisms
The privacy analysis of such mechanisms $\mathcal{M}$ is challenging as the creation of sub-mechanisms and their inputs is adaptive and dependent. The goal of this section is to give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be bounded by applying our concurrent composition theorems for continual mechanisms to the sub-mechanisms. Specifically, given that the $f^{\prime}$-concurrent composition of CMs is $(\epsilon,\delta)$-DP w.r.t. some verification function $f^{\prime}$, we would like to show that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the given verification function $f$.
@@
To prove that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. $f$, we need to show that the views of every adversary $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable. Intuitively, the messages that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ receives from $\mathcal{I}(b)$ contain information about $b$. Therefore, to deduce that $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f$ from the fact that ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$, the IPM $\mathcal{P}$ must not directly use the messages from $\mathcal{I}$ to respond. Instead, it should only use these messages to determine which queries it wants to send to its sub-mechanisms via ExtConComp. More precisely, $\mathcal{P}$ must generate an answer for $\mathcal{I}$ only based on the privatized answers provided by ExtConComp and independent of the un-privatized input sequence (Response Property). Moreover, $\mathcal{P}$ must decide whether to continue interacting with ExtConComp or to return a response based on these privatized answers (Destination Property).
@@
In Section 6.1, we formalize the above properties for $\mathcal{P}$, and in Section 6.2, we prove the desired privacy guarantees for $\mathcal{M}$. Combining this result with the concurrent composition theorems provided in the previous sections, we can avoid complicated privacy analysis for many CMs. To obtain rigorous results, our statements and proofs must be presented in a technical and somewhat complex form. In contrast, mechanisms are usually described more simply using pseudocode rather than message-passing state machines. In Section 6.3, we reformulate our results in terms of pseudocode, enabling new mechanisms to be analyzed directly without the need for reformulating to the setting in this paper.
@@
Recall that the goal of this section is to show $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f$ if ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$. That is, we need to show that the views of every adversary interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable.
The messages received by $\mathcal{P}$ from $\mathcal{I}$ contain information about the bit used by $\mathcal{I}$ to filter the adversary’s pairs. Therefore, to directly use the existing composition theorems for ExtConComp to analyze the privacy of $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$, we need to ensure that $\mathcal{P}$ only utilizes the messages from $\mathcal{I}$ to generate queries for the existing private sub-mechanisms executed by ExtConComp. Moreover, we need to guarantee that the sequence $(m_{1}^{0\prime},m_{1}^{1\prime}),(m_{2}^{0\prime},m_{2}^{1\prime}),\dots$ is $f^{\prime}$-valid, where $m_{1}^{b\prime},m_{2}^{b\prime},\dots$ is the sequence of messages that $\mathcal{P}$ sends to ExtConComp when it receives messages from $\mathcal{I}(b)\!\circ^{*}\!\mathcal{V}[f]$. The following properties formalize these:
@@
  • $f\to f^{\prime}$ Property. For $v\in\{{\texttt{L}},{\texttt{R}}\}$, let the pair $(m,v)$ denote a message $m$ sent to or received from the $v$’th mechanism (i.e., left or right) of $\mathcal{P}$. For $b\in\{0,1\}$, let $\sigma^{b}=\left((m_{1}^{b},v_{1}^{b}),\dots,(m_{t}^{b},v_{t}^{b})\right)$ be a sequence of $t$ input messages to $\mathcal{P}$, and let $\tau^{b}=\left((m_{1}^{b\prime},v_{1}^{b\prime}),\dots,(m_{t}^{b\prime},v_{t}^{b\prime})\right)$ be the corresponding sequence of $t$ output messages from $\mathcal{P}$. Suppose the following conditions hold:
@@
In Theorem 6.1, we assume that the views of every adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable. The goal is to show the same result for $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$. The key idea of the proof is to show that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{T}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical for some IPM $\mathcal{T}$ with a singleton initial state space. The proof consists of two steps.
@@
When $\mathcal{P}^{2}$ receives a message $m\in Q_{\mathcal{P}}^{L}$ from the left mechanism, it forwards two copies of $m$ to both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ as left messages. Specifically, it executes $(s_{0},v_{0}^{\prime},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{L}},m)$ and $(s_{1},v_{1}^{\prime},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{L}},m)$. When $\mathcal{P}^{2}$ receives a pair of messages $(m_{0},m_{1})\in Q_{\mathcal{P}}^{R}\times Q_{\mathcal{P}}^{R}$ from the right mechanism, it passes each $m_{b}$ to $\mathcal{P}_{b}$ as a right message. Specifically, it executes $(s_{0},v_{0}^{\prime},m_{0}^{\prime})\leftarrow\mathcal{P}_{0}(s_{0},{\texttt{R}},m_{0})$ and $(s_{1},v_{1}^{\prime},m_{1}^{\prime})\leftarrow\mathcal{P}_{1}(s_{1},{\texttt{R}},m_{0})$. By the destination property of $\mathcal{P}$ and the fact that both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ receive identical left messages, we have $v_{0}^{\prime}=v_{1}^{\prime}$. If $v_{0}^{\prime}=v_{1}^{\prime}={\texttt{R}}$, then $\mathcal{P}^{2}$ sends $(m_{0}^{\prime},m_{1}^{\prime})$ to its right mechanism. If $v_{0}^{\prime}=v_{1}^{\prime}={\texttt{L}}$, then by the response property, $m_{0}^{\prime}=m_{1}^{\prime}$, and $\mathcal{P}^{2}$ returns $m_{0}^{\prime}$ to its left mechanism.
The following claim follows directly from the definition of $\mathcal{P}^{2}$:
@@
$\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}$ in Claim 6.4 is an IPM with a singleton initial state space. Thus, by Lemma 2.16, the fact that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ are $(\epsilon,\delta)$-indistinguishable implies that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof. ∎